# Variational Diffusion Posterior Sampling with Midpoint Guidance

### v1

@@ B.2 Experimental setup
In the experiment described in Example 3.2, we set ${d}=100$ and generate $500$ instances of inverse problems $(\mathrm{A},m,\Sigma)$. For each instance, we compute the Wasserstein-2 distance between the resulting posterior distribution $\pi$ and $\smash{\hat{\pi}^{\ell(\eta)}_{0}}$.
@@
The mean of the prior is sampled from a standard Gaussian distribution. To generate the covariance $\Sigma\in\mathcal{S}^{++}_{d}$, we first draw a matrix $\mathrm{G}\in\mathbb{R}^{{d}\times{d}}$ with i.i.d. entries sampled from $\mathrm{N}(0,1)$. Then, for a better conditioning of $\Sigma$, we normalize the columns of $\mathrm{G}$ and set $\Sigma=\bar{\lambda}^{2}\mathrm{I}+\mathrm{G}\mathrm{G}^{\top}$, where $\bar{\lambda}^{2}$ is the mean of the squared singular values of $\mathrm{G}$.
@@
To sample an ill-posed problem we generate a rank-deficient matrix $\mathrm{A}\in\mathbb{R}^{{d}_{y}\times{d}}$ with ${d}_{y}\leq{d}$. We sample uniformly ${d}_{y}$ from the interval $\llbracket{d}/10,{d}\rrbracket$ and draw the entries of $\mathrm{A}$ i.i.d. from $\mathrm{N}(0,1)$. Regarding $\sigma_{y}$, we sample it uniformly from the interval $[0.1,0.5]$.
@@
$\pi(x)\propto g(x)p(x)\propto\mathcal{N}(x;m_{y},\Sigma_{y})\,,$
where $\Sigma_{y}=(\Sigma^{-1}+(1/\sigma_{y}^{2})\mathrm{A}^{\top}\mathrm{A})^{-1}$ and $m_{y}=\Sigma_{y}\big{(}(1/\sigma_{y}^{2})\mathrm{A}^{\top}y+\Sigma^{-1}m\big{)}$. The Wasserstein-2 distance between the true posterior and $\smash{\hat{\pi}^{\ell(\eta)}_{0}}$ is thus Olkin & Pukelsheim (1982)
$W_{2}(\pi,\smash{\hat{\pi}^{\ell}_{0}})^{2}=\|m_{y}-\hat{\mu}^{\ell}_{0}\|^{2}+\operatorname{tr}\big{(}\Sigma_{y}+\hat{\Sigma}^{\ell}_{0}-(\Sigma^{\frac{1}{2}}_{y}\hat{\Sigma}^{\ell}_{0}\Sigma^{\frac{1}{2}}_{y})^{\frac{1}{2}}\big{)}\,.$
@@ Appendix C More details on MGPS
C.1 Loss gradient
Recall that the loss function we optimize is $\varphi\mapsto\mathsf{KL}(\lambda^{\varphi}_{{\ell_{k}}|k+1}\parallel\smash{\hat{\pi}^{\theta}_{{\ell_{k}}|k+1}(\cdot|x_{k+1})})$ for a given $x_{k+1}$. Its gradient writes
$\nabla_{\varphi}\mathcal{L}_{k}(\varphi;X_{k+1})$ $=-\mathbb{E}\big{[}\nabla_{\varphi}\log\hat{g}^{\theta}_{{\ell_{k}}}(\hat{\mu}_{{\ell_{k}}}+\mathrm{diag}(\mathrm{e}^{\hat{\rho}_{\ell_{k}}})Z)\big{]}+\nabla_{\varphi}\mathsf{KL}(\lambda^{\varphi}_{{\ell_{k}}|k+1}\parallel p^{\theta}_{{\ell_{k}}|k+1}(\cdot|X_{k+1}))$
@@ C.1 Loss gradient
$\nabla_{\varphi}\mathsf{KL}(\lambda^{\varphi}_{{\ell_{k}}|k+1}\parallel p^{\theta}_{{\ell_{k}}|k+1}(\cdot|x_{k+1}))\\ =\nabla_{\varphi}\left[-\sum_{j=1}^{d}\hat{\rho}_{{\ell_{k}},j}+\frac{\|\hat{\mu}_{\ell_{k}}-m_{{\ell_{k}}|0,k+1}(\smash{m^{\theta}_{0|k+1}}(x_{k+1}),x_{k+1})\|^{2}+\sum_{j=1}^{d}\frac{\mathrm{e}^{2\hat{\rho}_{{\ell_{k}},j}}}{v_{{\ell_{k}}|0,k+1}}}{2v_{{\ell_{k}}|0,k+1}}\right]\,.$
@@ C.2 Warm start
In this section we describe the warm-start approach discussed in the main paper. The complete algorithm with the warm-start procedure is given in Algorithm 3. The original version presented in Algorithm 1 relies on first sampling approximately $X_{{\ell_{k}}}$, given $X_{k+1}$, from the surrogate transition $\smash{\hat{\pi}^{\theta}_{{\ell_{k}}|k+1}(\cdot|X_{k+1})}$ and then sampling $X_{k}$ from $q_{k|{\ell_{k}},k+1}(\cdot|X_{\ell_{k}},X_{k+1})$. In our warm-start approach, which we apply only during the first iterations of the algorithm (see the hyperparameter settings in Table 6), we draw inspiration from the decomposition
$\smash{\pi}_{k|k+1}(x_{k}|x_{k+1})=\int q_{k|1,k+1}(x_{k}|x_{1},x_{k+1})\smash{\pi}_{1|{\ell_{k}}}(x_{1}|x_{\ell_{k}})\smash{\pi}_{{\ell_{k}}|k+1}(x_{\ell_{k}}|x_{k+1})\,\mathrm{d}x_{1}\mathrm{d}x_{\ell_{k}}\,.$ (C.1)
@@
It suggests introducing a second intermediary step that involves sampling approximately from the transition $\smash{X_{1}\sim\smash{\pi}_{1|{\ell_{k}}}(\cdot|X_{\ell_{k}})}$. $X_{k}$ is then sampled from the bridge $q_{k|1,k+1}(\cdot|X_{1},X_{k+1})$.
@@
In order to draw approximate samples from $\smash{\pi}_{1|{\ell_{k}}}(\cdot|X_{{\ell_{k}}})$ we leverage again a Gaussian variational approximation $\smash{\lambda^{\psi}_{1|{\ell_{k}}}\vcentcolon=\mathrm{N}(\tilde{\mu}_{1},\mathrm{diag}(\mathrm{e}^{2\tilde{\rho}_{1}}))}$, which we fit by minimizing a proxy of the KL divergence between $\lambda^{\psi}_{1|{\ell_{k}}}$ and $\smash{\pi}_{1|{\ell_{k}}}(\cdot|X_{\ell_{k}})$ of which the gradient w.r.t. $\psi\vcentcolon=(\tilde{\mu}_{1},\tilde{\rho}_{1})$ writes
@@
$\nabla_{\psi}\mathcal{L}^{\mathrm{ws}}_{1|{\ell_{k}}}(\psi;X_{\ell_{k}})_{|\psi_{0}}\vcentcolon=-\nabla_{\psi}\left[\sum_{j=1}^{d}\tilde{\rho}_{1,j}-\frac{\|X_{\ell_{k}}-(\alpha_{{\ell_{k}}}/\alpha_{1})^{1/2}\tilde{\mu}_{1}\|^{2}+(\alpha_{{\ell_{k}}}/\alpha_{1})\sum_{j=1}^{d}\mathrm{e}^{\tilde{\rho}^{2}_{1,j}}}{2(1-\alpha_{{\ell_{k}}}/\alpha_{1})}\right]_{|\psi_{0}}\\ -\mathbb{E}\left[\nabla_{\psi}(\tilde{\mu}_{1}+\mathrm{diag}(\mathrm{e}^{\tilde{\rho}_{1}})Z)^{\intercal}_{|\psi_{0}}(\nabla_{x}\log g+s^{\theta}_{1})(\tilde{\mu}_{1}+\mathrm{diag}(\mathrm{e}^{\tilde{\rho}_{1}})Z\,{}_{|\psi_{0}})\right]\,.$ (C.2)
@@
This expression corresponds to the gradient w.r.t. $\psi$ of $\mathsf{KL}(\lambda^{\psi}_{1|{\ell_{k}}}\parallel\smash{\pi}_{1|{\ell_{k}}}(\cdot|X_{\ell_{k}}))$ combined with the score approximation $s^{\theta}_{1}$ of $\nabla_{x}\log q_{1}$ and the mild potential approximation $\nabla_{x}\log g_{1}$ of $\nabla_{x}\log g$. Indeed, we have that
@@
$\mathsf{KL}(\lambda^{\psi}_{1|{\ell_{k}}}\parallel\smash{\pi}_{1|{\ell_{k}}}(\cdot|X_{\ell_{k}}))=-\mathbb{E}_{\lambda^{\psi}_{1|{\ell_{k}}}}\big{[}\log g_{1}(X_{1})+\log q_{1}(X_{1})\big{]}\\ +\int\log\frac{\lambda^{\psi}_{1|{\ell_{k}}}(x_{1})}{q_{{\ell_{k}}|1}(X_{\ell_{k}}|x_{1})}\lambda^{\psi}_{1|{\ell_{k}}}(x_{1})\,\mathrm{d}x_{1}\,.$
@@
$\nabla_{\psi}$ $\mathbb{E}_{\lambda^{\varphi}_{1|{\ell_{k}}}}\big{[}\log g_{1}(X_{1})+\log q_{1}(X_{1})\big{]}_{|\psi_{0}}$
$=\nabla_{\psi}\mathbb{E}\left[\big{(}\log g_{1}+\log q_{1}\big{)}(\tilde{\mu}_{1}+\mathrm{diag}(\mathrm{e}^{\tilde{\rho}_{1}})Z)\right]_{|\psi_{0}}$
$=\mathbb{E}\left[\nabla_{\psi}(\tilde{\mu}_{1}+\mathrm{diag}(\mathrm{e}^{\tilde{\rho}_{1}})Z)^{\intercal}_{|\psi_{0}}(\nabla_{x}\log g_{1}+\nabla_{x}\log q_{1})(\tilde{\mu}_{1}+\mathrm{diag}(\mathrm{e}^{\tilde{\rho}_{1}})Z\,{}_{|\psi_{0}})\right]\,.$
Plugging the previous approximations yields the second term in (C.2). The warm-start procedure is summarized in Algorithm 2. We also use a single sample Monte Carlo estimate to perform the optimization. Finally, at step $k$ and given $\hat{X}_{\ell_{k}}$, the initial parameter $\psi_{k}$ of the variational approximation is chosen so that
$\lambda^{\psi_{k}}_{1|{\ell_{k}}}(x_{1})=q_{1|0,{\ell_{k}}}(x_{1}|\smash{m^{\theta}_{0|{\ell_{k}}}}(\hat{X}_{\ell_{k}}),\hat{X}_{\ell_{k}})\,.$
After having sampled $\tilde{X}_{1}\sim\lambda^{\psi^{*}_{k}}_{1|{\ell_{k}}}$, we use it to first sample $X_{k}\sim q_{k|1,k+1}(\cdot|\tilde{X}_{1},X_{k+1})$ and then initialize the next variational approximation $\lambda^{\varphi}_{{\ell_{k-1}}|k}$. The initial parameter $\varphi_{k-1}$ is set so that
$\lambda^{\varphi_{k-1}}_{{\ell_{k-1}}|k}(x_{\ell_{k-1}})=q_{{\ell_{k-1}}|0,k}(x_{\ell_{k-1}}|\tilde{X}_{1},X_{k})\,,$
@@
1:  Input: step $k$, samples $(\hat{X}_{\ell_{k}},X_{k+1})$, gradient steps $M$
2:  $\tilde{\mu}_{1}\leftarrow m_{1|0,{\ell_{k}}}(\smash{m^{\theta}_{0|{\ell_{k}}}}(\hat{X}_{\ell_{k}}),\hat{X}_{\ell_{k}})\,,\quad\tilde{\rho}_{1}\leftarrow\frac{1}{2}\log v_{1|0,{\ell_{k}}}$
@@
4:     $Z\sim\mathrm{N}(\mathrm{0}_{d},\mathrm{I}_{d})$
5:     $(\tilde{\mu}_{1},\tilde{\rho}_{1})\leftarrow\mathsf{OptimizerStep}(\nabla_{\psi}\widetilde{\mathcal{L}}^{\mathrm{ws}}_{1|{\ell_{k}}}(\cdot,Z;X_{\ell_{k}});\,\tilde{\mu}_{1},\tilde{\rho}_{1})$
@@
7:  $\hat{X}_{1}\leftarrow\hat{\mu}_{1}+\mathrm{diag}(\mathrm{e}^{\hat{\rho}_{1}})Z_{1}$ where $Z_{1}\sim\mathrm{N}(\mathrm{0}_{d},\mathrm{I}_{d})$
8:  $X_{k}\leftarrow m_{k|1,k+1}(\hat{X}_{1},X_{k+1})+v^{1/2}_{k|1,k+1}Z_{k}$
9:  Output: $X_{k},\hat{X}_{1}$
@@
2:  $X_{n}\sim\mathcal{N}(\mathrm{0}_{d},\mathrm{I}_{d}),\,\hat{X}_{n}\leftarrow X_{n}$
3:  $\hat{X}_{0}\leftarrow\smash{m^{\theta}_{0|n}}(\hat{X}_{n})$
@@
5:     $\hat{\mu}_{{\ell_{k}}}\leftarrow m_{{\ell_{k}}|0,k+1}(\hat{X}_{0},X_{k+1}),\quad\hat{\rho}_{\ell_{k}}\leftarrow\frac{1}{2}\log v_{{\ell_{k}}|0,k+1}$
@@
7:        $Z\sim\mathrm{N}(\mathrm{0}_{d},\mathrm{I}_{d})$
8:        $(\hat{\mu}_{\ell_{k}},\hat{\rho}_{\ell_{k}})\leftarrow\mathsf{OptimizerStep}(\nabla_{\varphi}\widetilde{\mathcal{L}}_{\ell_{k}}(\cdot,Z;X_{k+1});\,\hat{\mu}_{\ell_{k}},\hat{\rho}_{\ell_{k}})$

### v2

@@ B.2 Experimental setup
In the experiment described in Example 3.2, we set ${d}=100$ and generate $500$ instances of inverse problems $(\mathbf{A},\boldsymbol{m},\mathbf{\Sigma})$. For each instance, we compute the Wasserstein-2 distance between the resulting posterior distribution $\pi$ and $\smash{\hat{\pi}^{\ell(\eta)}_{0}}$.
@@
The mean of the prior is sampled from a standard Gaussian distribution. To generate the covariance $\mathbf{\Sigma}\in\mathcal{S}^{++}_{d}$, we first draw a matrix $\mathbf{G}\in\mathbb{R}^{{d}\times{d}}$ with i.i.d. entries sampled from $\mathrm{N}(0,1)$. Then, for a better conditioning of $\mathbf{\Sigma}$, we normalize the columns of $\mathbf{G}$ and set $\mathbf{\Sigma}=\bar{\lambda}^{2}\mathbf{I}+\mathbf{G}\mathbf{G}^{\top}$, where $\bar{\lambda}^{2}$ is the mean of the squared singular values of $\mathbf{G}$.
@@
To sample an ill-posed problem we generate a rank-deficient matrix $\mathbf{A}\in\mathbb{R}^{{d}_{{\boldsymbol{y}}}\times{d}}$ with ${d}_{{\boldsymbol{y}}}\leq{d}$. We sample uniformly ${d}_{{\boldsymbol{y}}}$ from the interval $\llbracket{d}/10,{d}\rrbracket$ and draw the entries of $\mathbf{A}$ i.i.d. from $\mathrm{N}(0,1)$. Regarding $\sigma_{\boldsymbol{y}}$, we sample it uniformly from the interval $[0.1,0.5]$.
@@
$\pi(\boldsymbol{x})\propto p({\boldsymbol{y}}|\boldsymbol{x})p(\boldsymbol{x})\propto\mathcal{N}(\boldsymbol{x};\boldsymbol{m}_{{\boldsymbol{y}}},\mathbf{\Sigma}_{{\boldsymbol{y}}})\,,$
where $\mathbf{\Sigma}_{{\boldsymbol{y}}}=(\mathbf{\Sigma}^{-1}+(1/\sigma_{\boldsymbol{y}}^{2})\mathbf{A}^{\top}\mathbf{A})^{-1}$ and $\boldsymbol{m}_{{\boldsymbol{y}}}=\mathbf{\Sigma}_{{\boldsymbol{y}}}\big{(}(1/\sigma_{\boldsymbol{y}}^{2})\mathbf{A}^{\top}y+\mathbf{\Sigma}^{-1}\boldsymbol{m}\big{)}$. The Wasserstein-2 distance between the true posterior and $\smash{\hat{\pi}^{\ell(\eta)}_{0}}$ is thus Olkin & Pukelsheim (1982)
$W_{2}(\pi,\smash{\hat{\pi}^{\ell}_{0}})^{2}=\|\boldsymbol{m}_{{\boldsymbol{y}}}-\hat{\boldsymbol{\mu}}^{\ell}_{0}\|^{2}+\operatorname{tr}\big{(}\mathbf{\Sigma}_{{\boldsymbol{y}}}+\hat{\mathbf{\Sigma}}^{\ell}_{0}-2(\mathbf{\Sigma}^{\frac{1}{2}}_{{\boldsymbol{y}}}\hat{\mathbf{\Sigma}}^{\ell}_{0}\mathbf{\Sigma}^{\frac{1}{2}}_{{\boldsymbol{y}}})^{\frac{1}{2}}\big{)}\,.$
@@ Appendix C More details on MGPS
C.1 Details on the loss
First, by the Data Processing inequality (Van Erven & Harremos, 2014, Example 2),
$\mathsf{KL}(\lambda^{{\boldsymbol{\varphi}}}_{k|k+1}(\cdot|\boldsymbol{x}_{k+1})\parallel\smash{\hat{\pi}^{\ell}_{k|k+1}(\cdot|\boldsymbol{x}_{k+1})})$
$\leq\int\log\frac{q_{k|{\ell_{k}},k+1}(\boldsymbol{x}_{k}|\boldsymbol{x}_{\ell_{k}},\boldsymbol{x}_{k+1})\lambda^{{\boldsymbol{\varphi}}}_{{\ell_{k}}|k+1}(\boldsymbol{x}_{\ell_{k}})}{q_{k|{\ell_{k}},k+1}(\boldsymbol{x}_{k}|\boldsymbol{x}_{\ell_{k}},\boldsymbol{x}_{k+1})\smash{\hat{\pi}^{\theta}_{{\ell_{k}}|k+1}(\boldsymbol{x}_{\ell_{k}}|\boldsymbol{x}_{k+1})}}q_{k|{\ell_{k}},k+1}(\boldsymbol{x}_{k}|\boldsymbol{x}_{\ell_{k}},\boldsymbol{x}_{k+1})\lambda^{{\boldsymbol{\varphi}}}_{{\ell_{k}}|k+1}(\boldsymbol{x}_{\ell_{k}})\,\mathrm{d}\boldsymbol{x}_{k}\mathrm{d}\boldsymbol{x}_{\ell_{k}}$
$=\int\log\frac{\lambda^{{\boldsymbol{\varphi}}}_{{\ell_{k}}|k+1}(\boldsymbol{x}_{\ell_{k}})}{\smash{\hat{\pi}^{\theta}_{{\ell_{k}}|k+1}(\boldsymbol{x}_{\ell_{k}}|\boldsymbol{x}_{k+1})}}q_{k|{\ell_{k}},k+1}(\boldsymbol{x}_{k}|\boldsymbol{x}_{\ell_{k}},\boldsymbol{x}_{k+1})\lambda^{{\boldsymbol{\varphi}}}_{{\ell_{k}}|k+1}(\boldsymbol{x}_{\ell_{k}})\,\mathrm{d}\boldsymbol{x}_{k}\mathrm{d}\boldsymbol{x}_{\ell_{k}}$
$=\mathsf{KL}(\lambda^{{\boldsymbol{\varphi}}}_{{\ell_{k}}|k+1}\parallel\smash{\hat{\pi}^{\theta}_{{\ell_{k}}|k+1}(\cdot|\boldsymbol{x}_{k+1})})$
The gradient of ${\boldsymbol{\varphi}}\mapsto\mathsf{KL}(\lambda^{{\boldsymbol{\varphi}}}_{{\ell_{k}}|k+1}\parallel\smash{\hat{\pi}^{\theta}_{{\ell_{k}}|k+1}(\cdot|\boldsymbol{x}_{k+1})})$ for a given $\boldsymbol{x}_{k+1}$ writes
$\nabla_{\boldsymbol{\varphi}}\mathcal{L}_{k}({\boldsymbol{\varphi}};\mathbf{X}_{k+1})\\ =-\mathbb{E}\big{[}\nabla_{\boldsymbol{\varphi}}\log\hat{p}^{\theta}_{{\ell_{k}}}({\boldsymbol{y}}|\hat{\boldsymbol{\mu}}_{{\ell_{k}}}+\mbox{diag}(\mathrm{e}^{\hat{\boldsymbol{\rho}}_{\ell_{k}}})\mathbf{Z})\big{]}+\nabla_{\boldsymbol{\varphi}}\mathsf{KL}(\lambda^{{\boldsymbol{\varphi}}}_{{\ell_{k}}|k+1}\parallel p^{\theta}_{{\ell_{k}}|k+1}(\cdot|\mathbf{X}_{k+1}))$ (C.1)
@@ C.1 Details on the loss
$\nabla_{\boldsymbol{\varphi}}\mathsf{KL}(\lambda^{{\boldsymbol{\varphi}}}_{{\ell_{k}}|k+1}\parallel p^{\theta}_{{\ell_{k}}|k+1}(\cdot|\boldsymbol{x}_{k+1}))\\ =\nabla_{\boldsymbol{\varphi}}\left[-\sum_{j=1}^{d}\hat{\boldsymbol{\rho}}_{{\ell_{k}},j}+\frac{\|\hat{\boldsymbol{\mu}}_{\ell_{k}}-\boldsymbol{m}_{{\ell_{k}}|0,k+1}(\smash{\boldsymbol{m}^{\theta}_{0|k+1}}(\boldsymbol{x}_{k+1}),\boldsymbol{x}_{k+1})\|^{2}+\sum_{j=1}^{d}\frac{\mathrm{e}^{2\hat{\boldsymbol{\rho}}_{{\ell_{k}},j}}}{v_{{\ell_{k}}|0,k+1}}}{2v_{{\ell_{k}}|0,k+1}}\right]\,.$
@@ C.2 Warm start
In this section we describe the warm-start approach discussed in the main paper. The complete algorithm with the warm-start procedure is given in Algorithm 3. The original version presented in Algorithm 1 relies on first sampling approximately $\mathbf{X}_{{\ell_{k}}}$, given $\mathbf{X}_{k+1}$, from the surrogate transition $\smash{\hat{\pi}^{\theta}_{{\ell_{k}}|k+1}(\cdot|\mathbf{X}_{k+1})}$ and then sampling $\mathbf{X}_{k}$ from $q_{k|{\ell_{k}},k+1}(\cdot|\mathbf{X}_{\ell_{k}},\mathbf{X}_{k+1})$. In our warm-start approach, which we apply only during the first iterations of the algorithm (see the hyperparameter settings in Table 6), we draw inspiration from the decomposition
$\smash{\pi}_{k|k+1}(\boldsymbol{x}_{k}|\boldsymbol{x}_{k+1})=\int q_{k|1,k+1}(\boldsymbol{x}_{k}|\boldsymbol{x}_{1},\boldsymbol{x}_{k+1})\smash{\pi}_{1|{\ell_{k}}}(\boldsymbol{x}_{1}|\boldsymbol{x}_{\ell_{k}})\smash{\pi}_{{\ell_{k}}|k+1}(\boldsymbol{x}_{\ell_{k}}|\boldsymbol{x}_{k+1})\,\mathrm{d}\boldsymbol{x}_{1}\mathrm{d}\boldsymbol{x}_{\ell_{k}}\,.$ (C.2)
@@
It suggests introducing a second intermediary step that involves sampling approximately from the transition $\smash{\mathbf{X}_{1}\sim\smash{\pi}_{1|{\ell_{k}}}(\cdot|\mathbf{X}_{\ell_{k}})}$. $\mathbf{X}_{k}$ is then sampled from the bridge $q_{k|1,k+1}(\cdot|\mathbf{X}_{1},\mathbf{X}_{k+1})$.
@@
In order to draw approximate samples from $\smash{\pi}_{1|{\ell_{k}}}(\cdot|\mathbf{X}_{{\ell_{k}}})$ we leverage again a Gaussian variational approximation $\smash{\lambda^{\psi}_{1|{\ell_{k}}}\vcentcolon=\mathrm{N}(\tilde{\boldsymbol{\mu}}_{1},\mbox{diag}(\mathrm{e}^{2\tilde{\boldsymbol{\rho}}_{1}}))}$, which we fit by minimizing a proxy of the KL divergence between $\lambda^{\psi}_{1|{\ell_{k}}}$ and $\smash{\pi}_{1|{\ell_{k}}}(\cdot|\mathbf{X}_{\ell_{k}})$ of which the gradient w.r.t. $\psi\vcentcolon=(\tilde{\boldsymbol{\mu}}_{1},\tilde{\boldsymbol{\rho}}_{1})$ writes
@@
$\nabla_{\psi}\mathcal{L}^{\mathrm{ws}}_{1|{\ell_{k}}}(\psi;\mathbf{X}_{\ell_{k}})_{|\psi_{0}}\vcentcolon=-\nabla_{\psi}\left[\sum_{j=1}^{d}\tilde{\boldsymbol{\rho}}_{1,j}-\frac{\|\mathbf{X}_{\ell_{k}}-(\alpha_{{\ell_{k}}}/\alpha_{1})^{1/2}\tilde{\boldsymbol{\mu}}_{1}\|^{2}+(\alpha_{{\ell_{k}}}/\alpha_{1})\sum_{j=1}^{d}\mathrm{e}^{\tilde{\boldsymbol{\rho}}^{2}_{1,j}}}{2(1-\alpha_{{\ell_{k}}}/\alpha_{1})}\right]_{|\psi_{0}}\\ -\mathbb{E}\left[\nabla_{\psi}(\tilde{\boldsymbol{\mu}}_{1}+\mbox{diag}(\mathrm{e}^{\tilde{\boldsymbol{\rho}}_{1}})\mathbf{Z})^{\intercal}_{|\psi_{0}}(\nabla_{x}\log p({\boldsymbol{y}}|\cdot)+\boldsymbol{s}^{\theta}_{1})(\tilde{\boldsymbol{\mu}}_{1}+\mbox{diag}(\mathrm{e}^{\tilde{\boldsymbol{\rho}}_{1}})\mathbf{Z}\,{}_{|\psi_{0}})\right]\,.$ (C.3)
@@
This expression corresponds to the gradient w.r.t. $\psi$ of $\mathsf{KL}(\lambda^{\psi}_{1|{\ell_{k}}}\parallel\smash{\pi}_{1|{\ell_{k}}}(\cdot|\mathbf{X}_{\ell_{k}}))$ combined with the score approximation $\boldsymbol{s}^{\theta}_{1}$ of $\nabla_{x}\log q_{1}$ and the mild likelihood approximation $\nabla_{x}\log p_{1}({\boldsymbol{y}}|\cdot)$ of $\nabla_{x}\log p({\boldsymbol{y}}|\cdot)$. Indeed, we have that
@@
$\mathsf{KL}(\lambda^{\psi}_{1|{\ell_{k}}}\parallel\smash{\pi}_{1|{\ell_{k}}}(\cdot|\mathbf{X}_{\ell_{k}}))=-\mathbb{E}_{\lambda^{\psi}_{1|{\ell_{k}}}}\big{[}\log p_{1}({\boldsymbol{y}}|\mathbf{X}_{1})+\log q_{1}(\mathbf{X}_{1})\big{]}\\ +\int\log\frac{\lambda^{\psi}_{1|{\ell_{k}}}(\boldsymbol{x}_{1})}{q_{{\ell_{k}}|1}(\mathbf{X}_{\ell_{k}}|\boldsymbol{x}_{1})}\lambda^{\psi}_{1|{\ell_{k}}}(\boldsymbol{x}_{1})\,\mathrm{d}\boldsymbol{x}_{1}\,.$
@@
$\nabla_{\psi}$ $\mathbb{E}_{\lambda^{{\boldsymbol{\varphi}}}_{1|{\ell_{k}}}}\big{[}\log p_{1}({\boldsymbol{y}}|\mathbf{X}_{1})+\log q_{1}(\mathbf{X}_{1})\big{]}_{|\psi_{0}}$
$=\nabla_{\psi}\mathbb{E}\left[\big{(}\log p_{1}({\boldsymbol{y}}|\cdot)+\log q_{1}\big{)}(\tilde{\boldsymbol{\mu}}_{1}+\mbox{diag}(\mathrm{e}^{\tilde{\boldsymbol{\rho}}_{1}})\mathbf{Z})\right]_{|\psi_{0}}$
$=\mathbb{E}\left[\nabla_{\psi}(\tilde{\boldsymbol{\mu}}_{1}+\mbox{diag}(\mathrm{e}^{\tilde{\boldsymbol{\rho}}_{1}})\mathbf{Z})^{\intercal}_{|\psi_{0}}(\nabla_{x}\log p_{1}({\boldsymbol{y}}|\cdot)+\nabla_{x}\log q_{1})(\tilde{\boldsymbol{\mu}}_{1}+\mbox{diag}(\mathrm{e}^{\tilde{\boldsymbol{\rho}}_{1}})\mathbf{Z}\,{}_{|\psi_{0}})\right]\,.$
Plugging the previous approximations yields the second term in (C.3). The warm-start procedure is summarized in Algorithm 2. We also use a single sample Monte Carlo estimate to perform the optimization. Finally, at step $k$ and given $\hat{\mathbf{X}}_{\ell_{k}}$, the initial parameter $\psi_{k}$ of the variational approximation is chosen so that
$\lambda^{\psi_{k}}_{1|{\ell_{k}}}(\boldsymbol{x}_{1})=q_{1|0,{\ell_{k}}}(\boldsymbol{x}_{1}|\smash{\boldsymbol{m}^{\theta}_{0|{\ell_{k}}}}(\hat{\mathbf{X}}_{\ell_{k}}),\hat{\mathbf{X}}_{\ell_{k}})\,.$
After having sampled $\tilde{\mathbf{X}}_{1}\sim\lambda^{\psi^{*}_{k}}_{1|{\ell_{k}}}$, we use it to first sample $\mathbf{X}_{k}\sim q_{k|1,k+1}(\cdot|\tilde{\mathbf{X}}_{1},\mathbf{X}_{k+1})$ and then initialize the next variational approximation $\lambda^{{\boldsymbol{\varphi}}}_{{\ell_{k-1}}|k}$. The initial parameter ${\boldsymbol{\varphi}}_{k-1}$ is set so that
$\lambda^{{\boldsymbol{\varphi}}_{k-1}}_{{\ell_{k-1}}|k}(\boldsymbol{x}_{\ell_{k-1}})=q_{{\ell_{k-1}}|0,k}(\boldsymbol{x}_{\ell_{k-1}}|\tilde{\mathbf{X}}_{1},\mathbf{X}_{k})\,,$
@@
1:  Input: step $k$, samples $(\hat{\mathbf{X}}_{\ell_{k}},\mathbf{X}_{k+1})$, gradient steps $M$
2:  $\tilde{\boldsymbol{\mu}}_{1}\leftarrow\boldsymbol{m}_{1|0,{\ell_{k}}}(\smash{\boldsymbol{m}^{\theta}_{0|{\ell_{k}}}}(\hat{\mathbf{X}}_{\ell_{k}}),\hat{\mathbf{X}}_{\ell_{k}})\,,\quad\tilde{\boldsymbol{\rho}}_{1}\leftarrow\frac{1}{2}\log v_{1|0,{\ell_{k}}}$
@@
4:     $\mathbf{Z}\sim\mathrm{N}(\mathbf{0}_{d},\mathbf{I}_{d})$
5:     $(\tilde{\boldsymbol{\mu}}_{1},\tilde{\boldsymbol{\rho}}_{1})\leftarrow\mathsf{OptimizerStep}(\nabla_{\psi}\widetilde{\mathcal{L}}^{\mathrm{ws}}_{1|{\ell_{k}}}(\cdot,\mathbf{Z};\mathbf{X}_{\ell_{k}});\,\tilde{\boldsymbol{\mu}}_{1},\tilde{\boldsymbol{\rho}}_{1})$
@@
7:  $\hat{\mathbf{X}}_{1}\leftarrow\hat{\boldsymbol{\mu}}_{1}+\mbox{diag}(\mathrm{e}^{\hat{\boldsymbol{\rho}}_{1}})\mathbf{Z}_{1}$ where $\mathbf{Z}_{1}\sim\mathrm{N}(\mathbf{0}_{d},\mathbf{I}_{d})$
8:  $\mathbf{X}_{k}\leftarrow\boldsymbol{m}_{k|1,k+1}(\hat{\mathbf{X}}_{1},\mathbf{X}_{k+1})+v^{1/2}_{k|1,k+1}\mathbf{Z}_{k}$
9:  Output: $\mathbf{X}_{k},\hat{\mathbf{X}}_{1}$
@@
2:  $\mathbf{X}_{n}\sim\mathcal{N}(\mathbf{0}_{d},\mathbf{I}_{d}),\,\hat{\mathbf{X}}_{n}\leftarrow\mathbf{X}_{n}$
3:  $\hat{\mathbf{X}}_{0}\leftarrow\smash{\boldsymbol{m}^{\theta}_{0|n}}(\hat{\mathbf{X}}_{n})$
@@
5:     $\hat{\boldsymbol{\mu}}_{{\ell_{k}}}\leftarrow\boldsymbol{m}_{{\ell_{k}}|0,k+1}(\hat{\mathbf{X}}_{0},\mathbf{X}_{k+1}),\quad\hat{\boldsymbol{\rho}}_{\ell_{k}}\leftarrow\frac{1}{2}\log v_{{\ell_{k}}|0,k+1}$
@@
7:        $\mathbf{Z}\sim\mathrm{N}(\mathbf{0}_{d},\mathbf{I}_{d})$
8:        $(\hat{\boldsymbol{\mu}}_{\ell_{k}},\hat{\boldsymbol{\rho}}_{\ell_{k}})\leftarrow\mathsf{OptimizerStep}(\nabla_{\boldsymbol{\varphi}}\widetilde{\mathcal{L}}_{\ell_{k}}(\cdot,\mathbf{Z};\mathbf{X}_{k+1});\,\hat{\boldsymbol{\mu}}_{\ell_{k}},\hat{\boldsymbol{\rho}}_{\ell_{k}})$